# Less is KEN: a Universal and Simple Non-Parametric Pruning Algorithm for Large Language Models

### v1

@@ Abstract
Neural network pruning has become increasingly crucial due to the complexity of neural network models and their widespread use in various fields. Existing pruning algorithms often suffer from limitations such as architecture specificity, excessive complexity and reliance on complex calculations, rendering them impractical for real-world applications. In this paper, we propose KEN: a straightforward, universal and unstructured pruning algorithm based on Kernel Density Estimation (KDE). KEN111Code available at https://github.com/itsmattei/KEN aims to construct optimized transformer models by selectively preserving the most significant parameters while restoring others to their pre-training state. This approach maintains model performance while allowing storage of only the optimized subnetwork, leading to significant memory savings. Extensive evaluations on seven transformer models demonstrate that KEN achieves equal or better performance than the original models with a minimum parameter reduction of 25%. In-depth comparisons against other pruning and PEFT algorithms confirm KEN effectiveness. Furthermore, we introduce KEN, an explainable tool that visualizes the optimized model composition and the subnetwork selected by KEN.
@@ 1 Introduction
Large Language Models (LLMs) have become the best and simplest solution for achieving state-of-the-art results in many natural language processing (NLP) applications. However, the increasing use of neural networks (NNs) and transformer models Vaswani et al. (2017) has resulted in a rise in computational cost due to the complexity of arithmetic calculations, larger matrices and the addition of more layers. Consequently, the weight and structure of these models become more complex, requiring high demands in computation and memory.
@@
One of the best approaches to address the overwhelming size of LLMs is to reduce their resources through pruning algorithms. These algorithms can eliminate parameters or entire components in a NN, making it lighter without compromising its original performance. Pruning algorithms emerged in parallel with the earliest use of NNs Mozer and Smolensky (1989); Janowsky (1989); LeCun et al. (1989), but they have gained significant importance in the last decade due to the widespread use of these networks in various fields. There are many pruning algorithms in literature Blalock et al. (2020), each with a unique approach or adapted old algorithms for these new architectures Benbaki et al. (2023). However, the complexity of neural networks can pose a challenge when creating pruning algorithms, as these may require creating new complex theories to make the models lightweight Dong et al. (2017); Malach et al. (2020). Additionally, existing pruning algorithms often exhibit shortcomings in their completeness Blalock et al. (2020) and fail to consider a critical aspect: the efficient storage of the pruned model. Some algorithms compress the model at runtime, but they lack a mechanism to persist the reduced NN for future use. Therefore, most algorithms in the literature focus only on the speed at which they reduce and execute the model without considering this crucial final stage. This is particularly important in resource-limited environments that use neural networks, such as smart devices and mobile phones Yang et al. (2017); Sze et al. (2017).
@@
This paper presents KEN (Kernel density Estimator for Neural network compression): a universal, simple, magnitude-based transformer pruning algorithm that leverages Kernel Density Estimation (KDE) for parameter pruning. In contrast to other pruning methods that rely on loss function minimization or exhaustive parameter search, KEN utilizes KDEs to identify and retain the most influential parameters while resetting the remaining ones to their original pre-trained values. This innovative pruning strategy streamlines the optimization process by leveraging the natural distribution of model parameters, eliminating any architecture-specific considerations. KEN effectively reduces the size of transformer models by a minimum of 25% without compromising performance. The pruned models consist solely of a subnetwork of trained parameters, which can be seamlessly downloaded and injected into pre-trained models on demand. This feature enables dynamic model reconfiguration and saves significant memory space that would otherwise be needed to store the fully trained model. Comparative evaluations demonstrate KEN exceptional capabilities, surpassing existing transformer pruning and PEFT algorithms. Additionally, we introduce KEN: an explainable tool that graphically depicts the optimized model from various perspectives. KEN highlights the KEN-selected parameters, their layer-wise differences and neighbor counts for each matrix that made up the analyzed model. Using KEN, we employed a non-parametric method widely used in statistics to create an efficient and intuitive pruning algorithm. Our approach achieved excellent results in terms of efficiency and performance, making it a practical alternative to other more complex pruning algorithms.
@@ 3 Related Work
(FLOP: Wang et al., 2019) FLOP is a magnitude-based pruning algorithm that employs matrix factorization to reduce the size of matrices in transformer models. This approach involves decomposing each matrix into smaller rank-1 components, which are then multiplied together to form the original matrix. For attention layers, FLOP decomposes each matrix into smaller rank-1 components based on the magnitudes of the matrix entries. For embedding layers, FLOP adaptively prunes dimensions based on word clusters. This means that FLOP only prunes dimensions that are not frequently used Joulin et al. (2017); Bastings et al. (2019), which helps to reduce the model size without sacrificing performance.
@@
(LoRA: Hu et al., 2021) LoRA is a novel fine-tuning method that leverages low-rank decomposition to reduce the parameter size of large language models (LLMs) while preserving their performance. This approach involves decomposing the LLMs weight matrices into low-rank components, which are then fine-tuned along with the original weights. This approach enables efficient parameter adaptation to specific tasks without compromising the LLMs generalization capabilities.
@@ 4 KEN pruning algorithm
KEN (Kernel density Estimator for Neural network compression) pruning algorithm is designed to identify and extract the most essential subnetwork from each transformer model following the main idea of the winning ticket hypothesis Frankle and Carbin (2018). This algorithm effectively prunes the network by employing Kernel Density Estimators (KDEs), retaining only the essential parameters and resetting the rest to their pre-trained values. The optimized subnetwork can be stored independently and seamlessly integrated into its pre-trained configuration for downstream applications.
KEN utilizes KDE to generalize the point distribution of each transformer matrix, resulting in a lightweight and smooth version of the original fine-tuned model. To prevent the complete deconstruction of the initial matrix composition, KEN applies KDE to the individual rows. The KDE calculation requires a value, which defines the number of points employed in the distribution calculation. Consequently, the value determines the number of the selected fine-tuned parameters, thus a lower value indicates a closer resemblance to the pre-trained model while a higher value reflects a closer alignment with its fine-tuned version.
The KEN algorithm can be described using the three phases defined below:
Phase 1: Parameter Extraction and KDE Calculation
Given a pre-trained matrix of a fixed layer :
@@
Phase 2: Parameter Retention and Pre-trained Value Reset
Using the KDE likelihood, the points that best fit the row distribution are identified, while the others are reset to their pre-trained values. This process results in an optimized row :
@@
Phase 3: Matrix Replacement and Optimized Fine-tuned Model
@@
KEN operates iteratively, replacing the matrix with during each iteration. So, after the t-th iteration, the model will have t-optimized matrices, effectively replacing the fine-tuned matrices without creating any additional versions of the model. This versatility allows KEN to prune the entire model or specific layer ranges.
Alg. 1 explains more formally all the three phases described to generate the optimized matrix . Additionally, the graphical representation displayed in Fig. 1 provides a clear and comprehensive visualization of all KEN steps while Fig.2 shows different matrices obtained using different values.
@@ 5.1 Experimental set-up
To evaluate KEN pruning algorithm performance across different architectures and datasets, we conducted a thorough series of experiments utilizing seven distinct transformer models. To maintain consistent evaluation conditions, we uniformly divided each dataset into training, validation and test sets. These divisions remained consistent throughout our experiments and across models. All datasets were imported from Huggingface222https://huggingface.co/datasets. To achieve optimal performance, we fine-tuned each model before applying KEN algorithm per each dataset, adjusting the number of epochs until the fine-tuned model achieved the highest possible F1-weighted score. Despite what the literature suggests, we used the F1 measure instead of classical accuracy as a comparison metric - if not explicitly used by the comparison benchmarks, because it delivers more reliable predictions, particularly on strongly unbalanced datasets.
To fully assess KEN algorithm capabilities, we gradually increased the value required for the algorithm, starting from a low value and incrementally increasing it until its fine-tuned version was reached. This incremental approach allowed us to identify the critical threshold value whereby the compressed model obtained results similar to its fine-tuned version or when the compression value leads to a catastrophic decline of performances, as reported in Apx. A.
@@
Model # Layers # params BLOOM Workshop et al. (2022) 24 1.72 B BLOOM Workshop et al. (2022) 24 560 M DeBERTa He et al. (2020) 12 138 M Bert Devlin et al. (2018) 12 109 M Ernie Sun et al. (2020) 12 109 M DistilBERT Sanh et al. (2019) 6 66 M Electra Clark et al. (2020) 12 33 M
Model Trainable params Reset params (%) AG-NEWS EMO IMDB YELP_POLARITY glue-sst2 BLOOM 442M 74.31 87.5 () 88.0 () 76.6 () 96.1 () 80.4 () 531M 69.17 92.2 () 90.6 () 84.2 () 96.3 () 90.9 () 664M 61.46 93.1 () 90.1 () 87.6 () 96.5 () 92.9 () BLOOM 411M 26.34 91.3 () 81.4 () 82.7 () 95.2 () 92.4 () 420M 24.80 91.8 () 83.0 () 84.3 () 95.3 () 92.1 () 429M 23.26 92.1 () 84.0 () 85.8 () 95.3 () 92.3 () DeBERTa 92M 33.86 92.2 () 87.9 () 82.5 () 95.9 () 94.6 () 99M 28.35 92.7 () 87.3 () 88.3 () 96.1 () 94.9 () 107M 22.84 92.9 () 87.1 () 89.8 () 96.2 () 94.8 () Bert 69M 37.05 93.4 () 84.2 () 86.8 () 95.0 () 93.7 () 75M 31.80 93.7 () 87.4 () 87.3 () 95.0 () 93.7 () 80M 26.55 93.6 () 87.9 () 87.6 () 95.1 () 93.8 () Ernie 69M 37.05 93.3 () 89.1 () 89.4 () 95.8 () 94.1 () 75M 31.80 93.3 () 88.7 () 89.2 () 95.8 () 93.8 () 80M 26.55 93.8 () 88.1 () 89.6 () 95.9 () 93.4 () DistilBERT 44M 34.39 92.3 () 88.1 () 83.2 () 94.6 () 91.9 () 47M 28.92 93.1 () 88.8 () 84.4 () 94.7 () 91.9 () 51M 23.45 93.3 () 88.2 () 84.6 () 94.9 () 92.0 () Electra 8.9M 75.56 84.1 () 84.3 () 78.9 () 88.5 () 79.9 () 12M 64.75 89.7 () 86.0 () 82.0 () 92.1 () 85.0 () 14M 55.94 91.3 () 85.6 () 84.3 () 93.7 () 90.1 ()
@@ 5.2 Model compression
Transformer models and other neural networks often have large file sizes, with a fine-tuned transformer potentially reaching up from 500 MB to 2GB or more in size. However, the KEN algorithm reduces this size by selecting and retaining a subset of parameters while restoring the rest to their pre-trained values. This process creates a more concentrated model that only includes the essential values for each matrix, resulting in significant weight reduction. To accurately assess the weight reduction achieved by KEN, we save the compressed model generated during this phase and compare it to its original, unpruned version. To ensure a fair comparison, we use the same technique to save both the compressed and original fine-tuned models. However, KEN requires a support file, such as a dictionary, to load the parameters saved into their appropriate positions during the loading process. This is because during loading, the fine-tuning values must be loaded into a pre-trained model and the support file provides the necessary mapping to ensure proper placement. Sec. 6.2 provides a comprehensive overview of the compression results obtained during this analysis.
@@ 6 Results and Discussion
In this section, we present the results obtained for each KEN main goals. Sec. 6.1 discusses the effectiveness of KEN-pruned models in comparison to: their unpruned counterparts, pruning benchmarks and state-of-the-art PEFT algorithm. Sec. 6.2 focuses on the process of saving and loading the subnetwork extracted by KEN, comparing the reduced file sizes achieved by it with those of the original models. Finally, Sec. 6.3 shows KEN, illustrating its applications.
@@ 6.1 Experiment results
To evaluate the efficacy of KEN, we conducted a series of experiments across diverse classification and sentiment analysis datasets. For each dataset, we implemented KEN multiple times, employing varied values and calculating the mean and standard deviation of the resulting F1-weighted scores. The complete dataset list can be found in Apx. B. As evidenced in Tab. 2, KEN successfully compressed all analyzed models without sacrificing their original, unpruned performance. We observed a remarkable reduction in overall model parameter count, ranging from a minimum of 25% to a substantial 70% for certain models. Intriguingly, the models with both the highest and lowest parameter counts exhibited the most significant parameter reduction. Additionally, for each model under examination, we observed no substantial difference in performance as the percentage of reset parameters increased, maintaining a remarkable resemblance to the unpruned model performance. This observation underscores KEN exceptional generalization capability, striking a balance between performance and compression even at middle-to-high compression rates.
We compared KEN to other pruning algorithms specifically designed for transformer models, including FLOP, Hybrid and HybridNT, as described in Sec. 3. It is essential to note that Lagunas et al. (2021) models (Hybrid and HybridNT) only prune the attention layers and not the entire model. To facilitate a comprehensive and standardized comparison of all algorithms, we recalibrated the size of their models based on our holistic perspective, ignoring any partial considerations. We combined the results obtained in their publication with those obtained from KEN and FLOP in Tab. 3. KEN outperformed all other compared models with a significant performance gap while utilizing fewer parameters in every instance. In addition to these findings, we conducted a thorough analysis of FLOP, which is the most complete pruning algorithm studied and, like KEN, decomposes original matrices to derive pruned ones. We conducted additional experiments on all models where FLOP could be applied, using the datasets listed in Tab. 2. We compared the results obtained from FLOP with those of KEN, which employed fewer parameters than FLOP. As shown in Tab. 4, FLOP outperforms KEN in only one instance. For all other models and datasets analyzed, KEN consistently outperforms FLOP.
@@
Model Pruning algorithm Trainable params. AG-NEWS EMO IMDB YELP_POLARITY glue-sst2 BLOOM KEN 531M 92.2 () 90.6 () 84.2 () 96.3 () 90.9 () FLOP 1.1B 90.1 () 84.0 () 80.9 () 85.5 () 80.7 () BLOOM KEN 404M 91.3 () 85.5 () 81.3 () 94.8 () 92.0 () FLOP 408M 91.0 () 84.0 () 72.1 () 87.0 () 81.8 () DeBERTa KEN 84M 91.4 () 88.9 () 82.5 () 96.0 () 92.8 () FLOP 88M 90.6 () 83.1 () 81.1 () 91.4 () 82.3 () Bert KEN 57M 91.6 () 86.0 () 84.9 () 93.8 () 92.8 () FLOP 66M 90.9 () 83.3 () 80.5 () 90.2 () 83.2 () Ernie KEN 57M 91.5 () 88.3 () 87.6 () 95.7 () 94.1 () FLOP 67M 89.8 ( ) 83.8 () 81.1 () 90.9 () 83.2 () DistilBERT KEN 40M 91.9 () 88.2 () 78.1 () 94.1 () 89.2 () FLOP 45M 90.7 () 83.2 () 81.2 () 90.7 () 82.4 () Electra KEN 14M 91.3 () 85.6 () 84.3 () 93.7 () 90.1 () FLOP 28M 90.9 () 83.1 () 81.2 () 90.5 () 81.1 ()
Although KEN belongs to the winning ticket pruning algorithms family, it shares similarities with Parameter Efficient Fine-tuning (PEFT) algorithms. This is because both approaches aim to identify a subset of optimal parameters within the fine-tuned model. We conducted a thorough evaluation of KEN and compared it to LoRA, which is currently the state-of-the-art PEFT algorithm. We applied LoRA and KEN to the same layers of each model. We then trained the LoRA-based models for five times more epochs than their KEN-based counterparts. Additionally, we gradually increased the number of rank decomposition matrices for each model from 16 to 768, which is the average size of the matrices in the tested models. In each LoRA-based experiment, only the LoRA-specific parameters were designated as either trainable or not. Our results, presented in Fig. 3, demonstrate that KEN consistently outperforms LoRA in terms of F1-measure while utilizing fewer trained parameters. However, when LoRA parameters are not the only ones trained, KEN and LoRA generally produce similar results. It is worth noting that LoRA consistently employs a larger parameter count than KEN. These compelling results provide strong evidence supporting our hypothesis that strategically selecting a subset of parameters and resetting the remainder offers a promising alternative to conventional pruning techniques.

### v2

@@ Abstract
Neural network pruning has become increasingly crucial due to the complexity of these models and their widespread use in various fields. Existing pruning algorithms often suffer from limitations such as architecture specificity, excessive complexity and reliance on demanding calculations, rendering them impractical for real-world applications. This paper introduces KEN: a straightforward, universal and unstructured pruning algorithm based on Kernel Density Estimation (KDE). KEN111Code available at https://github.com/itsmattei/KEN aims to construct optimized transformers by selectively preserving the most significant parameters while restoring others to their pre-training state. This strategy preserves model performance while enabling storage of only the optimized subnetwork, leading to substantial memory savings. Extensive evaluations across seven different LLMs demonstrate that KEN achieves equal or better performance than their original unpruned versions, with a minimum parameter reduction of 25%. Furthermore, in-depth comparisons with established pruning and PEFT algorithms confirm KEN effectiveness. We further introduce KENviz, an explainable tool that visualizes the optimized model composition achieved by KEN from different points of view.
@@ 1 Introduction
Large Language Models (LLMs) have become the best and simplest solution for achieving state-of-the-art results in many natural language processing (NLP) applications. However, the increasing use of neural networks (NNs) and transformers Vaswani et al. (2017) has resulted in a rise in computational cost due to the complexity of arithmetic calculations, larger matrices and the addition of more layers. Consequently, the weight and structure of these models become more complex, requiring high demands in computation and memory.
@@
One of the best approaches to address the overwhelming size of LLMs is to reduce their resources through pruning algorithms. These algorithms can eliminate parameters or entire components in a NN, making it lighter without compromising its original performance. Pruning algorithms emerged in parallel with the earliest use of NNs Mozer and Smolensky (1989); Janowsky (1989); LeCun et al. (1989), but they have gained significant importance in the last decade due to the widespread use of these networks in various fields. There are many pruning algorithms in literature Blalock et al. (2020), each with a unique approach or adapted old algorithms for these new architectures Benbaki et al. (2023). However, the complexity of neural networks can pose a challenge when creating pruning algorithms, as these may require new complex theorems to make the models lightweight Dong et al. (2017); Malach et al. (2020). Additionally, existing pruning algorithms often exhibit shortcomings in their completeness Blalock et al. (2020) and fail to consider a critical aspect: the efficient storage of the pruned result. Some algorithms compress models at runtime but lack mechanisms to preserve the reduced NN for future use. Consequently, most algorithms prioritize the speed of reduction and execution, neglecting this critical final stage essential in resource-limited environments Yang et al. (2017); Sze et al. (2017).
@@
This paper presents KEN (Kernel density Estimator for Neural network compression): a universal, simple, magnitude-based transformer pruning algorithm that leverages Kernel Density Estimation (KDE) for parameter pruning. Unlike other pruning methods that rely on minimizing loss functions or exhaustive parameter search, KEN - inspired by the winning ticket pruning hypothesis Frankle and Carbin (2018) - identifies and retains the most influential parameters using KDEs while resetting the others to their original pre-trained values. This innovative approach streamlines the optimization process by leveraging the natural distribution of model parameters, eliminating any architecture-specific considerations. KEN effectively reduces the size of transformer models by a minimum of 25% without compromising performance. The pruned models consist only of a subnetwork of trained parameters, which can be seamlessly downloaded and injected into its pre-trained version as needed. This feature enables dynamic model reconfiguration and saves significant memory space that would otherwise be required to store the fully trained model. Comparative evaluations demonstrate KEN exceptional capabilities, surpassing existing transformer pruning and PEFT algorithms. Finally, we introduce KENviz: an explainable tool that graphically depicts the optimized model from various perspectives. KENviz highlights the KEN-selected parameters, their layer-wise differences and neighbor counts for each matrix that made up the analyzed model. Using KEN, we employed a non-parametric method widely used in statistics, to create an efficient and intuitive pruning algorithm. Our approach achieved excellent results in terms of efficiency and performance, making it a practical alternative to other more complex pruning algorithms.
@@ 3 Related Work
(FLOP: Wang et al., 2019) is a magnitude-based pruning algorithm that employs matrix factorization to reduce the size of matrices in transformer models. This approach involves decomposing each matrix into rank components, which are then multiplied together to form the original matrix. For attention layers, FLOP decomposes each matrix into smaller rank-1 components based on the magnitudes of the matrix entries. Instead, for embedding layers, FLOP adaptively prunes dimensions based on word clusters. This means that FLOP only prunes dimensions that are not frequently used Joulin et al. (2017); Bastings et al. (2019), which helps to reduce the model size without sacrificing performance.
@@
(LoRA: Hu et al., 2021) is a novel fine-tuning method that leverages low-rank decomposition to reduce the parameter size of LLMs while preserving their performance. This approach involves decomposing LLMs weight matrices into low-rank components, which are then fine-tuned along with the original weights. LoRA enables efficient parameter adaptation to specific tasks without compromising the model generalization capabilities.
@@ 4 KEN pruning algorithm
KEN (Kernel density Estimator for Neural network compression) pruning algorithm is designed to identify and extract the most essential subnetwork from transformer models following the main idea of the winning ticket hypothesis Frankle and Carbin (2018). Our algorithm effectively prunes the network by employing Kernel Density Estimators (KDEs), retaining only the essential parameters and resetting the rest to their pre-trained values. The optimized subnetwork can be stored independently and seamlessly integrated into its pre-trained configuration for downstream applications.
KEN through KDEs, generalizes the point distribution of each transformer matrix, capturing the smoothed version of the original fine-tuned model. To prevent the complete deconstruction of the initial matrix composition, KEN applies KDEs to individual rows. The KDE calculation requires a value, which defines the number of points employed in the distribution calculation which directly influences the number of retained fine-tuned parameters (Fig. 1). Thus, a lower value indicates a closer resemblance to the pre-trained model while a higher value reflects a closer alignment with its fine-tuned version.
KEN algorithm operates in three primary steps:
Step 1: Parameter Extraction and KDE Calculation
Given a pre-trained matrix of a layer :
@@
Step 2: Parameter Retention and Pre-trained Value Reset
The points that best fit the row distribution are identified using the KDE likelihood, while the others are reset to their pre-trained values. This process results in an optimized row :
@@
Step 3: Matrix Replacement and Optimized Fine-tuned Model
@@
KEN operates iteratively, replacing the matrix with during each iteration. Therefore, after the iteration, the model will have matrices, effectively replacing the fine-tuned matrices without creating any additional versions of the model. This versatility allows KEN to prune the entire model or specific layer ranges.
Algorithm 1 provides a more formal explanation of the three steps described for generating the optimized matrix . Additionally, the graphical representation in Fig. 2 offers a clear and comprehensive visualization of all KEN steps, while Fig. 3 displays different matrices obtained using various values.
@@ 5.1 Experimental set-up
To evaluate KEN pruning algorithm performance across different architectures and datasets, we conducted a thorough series of experiments using seven distinct transformer models. To maintain consistent evaluation conditions, we uniformly divided each dataset into training, validation and test sets. These divisions remained consistent throughout our experiments and across models. All datasets were imported from Huggingface222https://huggingface.co/datasets. To achieve optimal performance, we fine-tuned each model before applying KEN algorithm per each dataset, adjusting the number of epochs until the fine-tuned model achieved the best F1-weighted score. Despite what the literature suggests, we used the F1 measure instead of classical accuracy as a comparison metric - if not explicitly used by the comparison benchmarks - because it delivers more reliable predictions, particularly on strongly unbalanced datasets.
To fully assess KEN capabilities, we gradually increased the value required by the algorithm, starting from a low value and incrementally increasing it until its fine-tuned version was reached. This incremental approach allowed us to identify the critical threshold value whereby the compressed model obtained results similar to its fine-tuned version or when the compression value leads to a catastrophic decline of performances, as reported in Apx. A.
@@
Model # Layers # params BLOOM1B7 Workshop et al. (2022) 24 1.72 B BLOOM560k Workshop et al. (2022) 24 560 M DeBERTa He et al. (2020) 12 138 M Bert Devlin et al. (2018) 12 109 M Ernie Sun et al. (2020) 12 109 M DistilBERT Sanh et al. (2019) 6 66 M Electra Clark et al. (2020) 12 33 M
Model Trainable params Reset params (%) AG-NEWS EMO IMDB YELP_POLARITY glue-sst2 BLOOM1B7 442M 74.31 87.5 () 88.0 () 76.6 () 96.1 () 80.4 () 531M 69.17 92.2 () 90.6 () 84.2 () 96.3 () 90.9 () 664M 61.46 93.1 () 90.1 () 87.6 () 96.5 () 92.9 () BLOOM560k 411M 26.34 91.3 () 81.4 () 82.7 () 95.2 () 92.4 () 420M 24.80 91.8 () 83.0 () 84.3 () 95.3 () 92.1 () 429M 23.26 92.1 () 84.0 () 85.8 () 95.3 () 92.3 () DeBERTa 92M 33.86 92.2 () 87.9 () 82.5 () 95.9 () 94.6 () 99M 28.35 92.7 () 87.3 () 88.3 () 96.1 () 94.9 () 107M 22.84 92.9 () 87.1 () 89.8 () 96.2 () 94.8 () Bert 69M 37.05 93.4 () 84.2 () 86.8 () 95.0 () 93.7 () 75M 31.80 93.7 () 87.4 () 87.3 () 95.0 () 93.7 () 80M 26.55 93.6 () 87.9 () 87.6 () 95.1 () 93.8 () Ernie 69M 37.05 93.3 () 89.1 () 89.4 () 95.8 () 94.1 () 75M 31.80 93.3 () 88.7 () 89.2 () 95.8 () 93.8 () 80M 26.55 93.8 () 88.1 () 89.6 () 95.9 () 93.4 () DistilBERT 44M 34.39 92.3 () 88.1 () 83.2 () 94.6 () 91.9 () 47M 28.92 93.1 () 88.8 () 84.4 () 94.7 () 91.9 () 51M 23.45 93.3 () 88.2 () 84.6 () 94.9 () 92.0 () Electra 8.9M 75.56 84.1 () 84.3 () 78.9 () 88.5 () 79.9 () 12M 64.75 89.7 () 86.0 () 82.0 () 92.1 () 85.0 () 14M 55.94 91.3 () 85.6 () 84.3 () 93.7 () 90.1 ()
@@ 5.2 Model compression
Transformer models, like many neural networks, often have large file sizes. A fine-tuned transformer can range from 500 MB to 2GB or more. However, the KEN algorithm reduces this size by selecting and retaining a subset of parameters while restoring the rest to their pre-trained values. This process creates a more concentrated model that only includes the essential values for each matrix, resulting in significant weight reduction. To quantify the effectiveness of KEN, we save the compressed model generated during this phase and compare it to its original, unpruned version. To ensure a fair comparison, we use the same technique to save both the compressed and original fine-tuned models. However, KEN requires a support file, such as a dictionary, to load the parameters saved into their appropriate positions during the loading process. This is because during loading, the fine-tuning values must be loaded into a pre-trained model and the support file provides the necessary mapping to ensure proper placement. Sec. 6.2 provides a comprehensive overview of the compression results obtained during this analysis.
@@ 6 Results and Discussion
In this section, we present the results obtained for each KEN main goals. Sec. 6.1 discusses the effectiveness of KEN-pruned models compared to their unpruned counterparts, pruning benchmarks and the state-of-the-art PEFT algorithm. Sec. 6.2 focuses on the process of saving and loading the subnetwork extracted by KEN, comparing the reduced file sizes achieved by it with those of the original models. Finally, Sec. 6.3 shows KENviz, illustrating its applications.
@@ 6.1 Experiment results
To evaluate the efficacy of KEN, we conducted a series of experiments across different classification and sentiment analysis datasets. For each dataset, we implemented KEN multiple times, employing varied values and calculating the mean and standard deviation of the resulting F1-weighted scores. The complete dataset list can be found in Apx. B. As evidenced in Tab. 2, KEN successfully compressed all analyzed models without sacrificing their original, unpruned performance. We observed a remarkable reduction in overall model parameter count, ranging from a minimum of 25% to a substantial 70% for certain models. Intriguingly, the models with both the highest and lowest parameter counts exhibited the most significant parameter reduction. Additionally, for each model under examination, we observed no substantial difference in performance as the percentage of reset parameters increased, maintaining a remarkable resemblance to the unpruned model performance. This observation underscores KEN exceptional generalization capability, balancing performance and compression even at middle-high compression rates.
We compared KEN to other pruning algorithms specifically designed for transformer models, including FLOP, Hybrid and HybridNT as described in Sec. 3. It is essential to note that Lagunas et al. (2021) models (Hybrid and HybridNT) only prune the attention layers and not the entire model. To facilitate a comprehensive and standardized comparison of all algorithms, we recalibrated the size of their models based on our holistic perspective, ignoring any partial considerations. We combined the results obtained in their publication with those obtained from KEN and FLOP in Tab. 3. KEN outperformed all other compared models with a significant performance gap while utilizing fewer parameters in every instance. In addition to these findings, we conducted a thorough analysis of FLOP, which is the most complete pruning algorithm studied and, like KEN, decomposes original matrices to derive pruned ones. We conducted additional experiments on all models where FLOP could be applied, using the datasets listed in Tab. 2. We compared the results obtained from FLOP with those of KEN, which employed fewer parameters than FLOP. As shown in Tab. 4, FLOP outperforms KEN in only one instance. For all other models and datasets analyzed, KEN consistently outperforms FLOP.
@@
Model Pruning algorithm Trainable params. AG-NEWS EMO IMDB YELP_POLARITY glue-sst2 BLOOM1B7 KEN 531M 92.2 () 90.6 () 84.2 () 96.3 () 90.9 () FLOP 1.1B 90.1 () 84.0 () 80.9 () 85.5 () 80.7 () BLOOM560k KEN 404M 91.3 () 85.5 () 81.3 () 94.8 () 92.0 () FLOP 408M 91.0 () 84.0 () 72.1 () 87.0 () 81.8 () DeBERTa KEN 84M 91.4 () 88.9 () 82.5 () 96.0 () 92.8 () FLOP 88M 90.6 () 83.1 () 81.1 () 91.4 () 82.3 () Bert KEN 57M 91.6 () 86.0 () 84.9 () 93.8 () 92.8 () FLOP 66M 90.9 () 83.3 () 80.5 () 90.2 () 83.2 () Ernie KEN 57M 91.5 () 88.3 () 87.6 () 95.7 () 94.1 () FLOP 67M 89.8 ( ) 83.8 () 81.1 () 90.9 () 83.2 () DistilBERT KEN 40M 91.9 () 88.2 () 78.1 () 94.1 () 89.2 () FLOP 45M 90.7 () 83.2 () 81.2 () 90.7 () 82.4 () Electra KEN 14M 91.3 () 85.6 () 84.3 () 93.7 () 90.1 () FLOP 28M 90.9 () 83.1 () 81.2 () 90.5 () 81.1 ()
Although KEN belongs to the winning ticket pruning algorithms family, it shares similarities with Parameter Efficient Fine-tuning (PEFT) algorithms. This is because both approaches aim to identify a subset of optimal parameters within the fine-tuned model. We thoroughly evaluated KEN and compared it to LoRA, which is currently the state-of-the-art PEFT algorithm. We applied LoRA and KEN to the same layers of each model and then trained the LoRA-based models using five times more training epochs than their KEN-based counterparts. Additionally, we gradually increased the number of rank decomposition matrices for each model from 16 to 768, which is the average size of the matrices in the tested models. In each LoRA-based experiment, only the LoRA-specific parameters were designated as either trainable or not. Our results, presented in Fig. 4, demonstrate that KEN consistently outperforms LoRA in terms of F1-measure while utilizing fewer trained parameters. However, when LoRA parameters are not the only ones trained, KEN and LoRA generally produce similar results. Nevertheless, LoRA consistently requires a larger parameter count than KEN. These compelling results provide strong evidence supporting our hypothesis that strategically selecting a subset of parameters and resetting the remainder offers a promising alternative to conventional pruning techniques.